OpenStudy (anonymous):

write the equation in slope-intercept form. y+3=3(x-1) & y-5=3/4(x-5) slope intercept form is y=mx+b m=slope b=y-intercept.

OpenStudy (anonymous):

slope intercept form is y=mx+b all you have to do is isolate y on the left side of the = sign, by subtracting the 3 on the left side and doing the same to the right side of the =. then expand the expressions on the right side of the = and combine like terms and put them in y=ms+b format.

OpenStudy (anonymous):

sorry for typo it should be y=mx+b